OpenStudy (anonymous):

simplfy sin(arccos0 + arccos1/2)

OpenStudy (loser66):

cos (pi/2) =?

OpenStudy (flvs.net):

Still need help?

OpenStudy (anonymous):

simplify [0, 180º] sin(arccos 0 + arccos 1/2) .. arccos 0=90º arccos 1/2=60º .. sin(arccos 0 + arccos 1/2) sin(90º+ 60º) =sin 150º =1/2

OpenStudy (anonymous):

did tht help?

OpenStudy (anonymous):

To approach it from a slightly different direction: Let \(x=\arccos0\) (so that \(\cos x=0\)) and \(y=\arccos\dfrac{1}{2}\) (so that \(\cos y=\dfrac{1}{2}\)). \[\begin{align*}\sin(x+y)&=\sin x\cos y+\cos x\sin y\\[1ex] &=\frac{1}{2}\sin x \end{align*}\] Since \(\cos x=0\), we have that \(x=\dfrac{k\pi}{2}\) for nonzero integer \(k\). Keeping in mind that \(0\le\arccos z\le\pi\), we omit all values of \(k\) except \(k=1\), so \(x=\dfrac{\pi}{2}\). Finally, what's \(\dfrac{1}{2}\sin\dfrac{\pi}{2}\)?